write an equation for the line that contains the point (-1 2) and is perpendicular to the line 3x+5y=7
Answer by robertb(5830)

The line would have the form 5x - 3y = c, for some particular constant c. To find this value, substitute the coordinates of the point into this equation:
5*(-1) - 3*2 = -5 - 6 = -11 = c.
Therefore, the line that contains the point (-1 2) and is perpendicular to the line 3x+5y=7 is 5x - 3y = -11.